A normal distribution has a mean of 80 and a standard deviation of 14.

Determine the value above which 80% of the values will occur. (Round z-score computation and your final answer to 2 decimal places.

Answer by stanbon(75887)

: A normal distribution has a mean of 80 and a standard deviation of 14.
Determine the value above which 80% of the values will occur. (Round z-score computation and your final answer to 2 decimal places.
--------
Find the z-score with a left tail of 0.20
invNorm(0.20) = -0.8416
-------
Ans: x = z*s + u
x = -0.8416*14 + 80 = 68.22
-------------------
Cheers,
Stan H.
-----------------